OpenStudy (anonymous):

If a line crosses the y-axis at (0,1) and has a slope of 4/5, what is the equation of the line? A) 4y-5x=5 B) y-4x=5 C) 5y+4x=5 D) 5y-4x=5 Thank you. Xoxoxo<3

OpenStudy (anonymous):

d

OpenStudy (ybarrap):

 Rearrange y=mx+b $$ y=(4/5)x+1 $$ Where 1 is the y-intercept, b, which we were given: (0,1) and (4/5) is m, the slope. $$ y=\cfrac{4x}{5}+1\\ y-\cfrac{4x}{5}=1\\ 5y-4x=5 $$ Does this make sense?

OpenStudy (welshfella):

y = mx + b is the general form of the slope-intercept equation m = slope and b = intercept on y-axis s here m = 4/5 and b = 1 y = (4/5)x + 1

OpenStudy (anonymous):

Thank you!! I wish I was good at math ! :(

OpenStudy (welshfella):

multiply through by 5 and rearrange to get the correct option